OpenStudy (watchmath):

What is the minimum value of \(|\sin x+\cos x+\tan x+\cot x+\sec x+\csc x|\)

OpenStudy (faiqraees):

Start by finding the derivative of the function. Equate the derivative to zero and then obtain a value for x. Plug the value of x in the given equation and you will have the minimum value

OpenStudy (faiqraees):

\[\rm \frac{d}{dx} |f(x)| = \frac{f(x)}{|f(x)|} \times \frac{d}{dx} f(x)\] \[\rm \frac{d}{dx} \sin x = \cos x\]\[\rm \frac{d}{dx} \cos x = -\sin x\]\[\rm \frac{d}{dx} \tan x = \sec^2 x\]\[\rm \frac{d}{dx} \sec x = \sec x \tan x\]\[\rm \frac{d}{dx} \cot x = \csc^2 x\]\[\rm \frac{d}{dx} \csc x = -\cot x \csc x \]

OpenStudy (watchmath):

Let see if you manage to find the stationary point(s)

OpenStudy (eliesaab):

The minimum is about 1.82843

OpenStudy (eliesaab):

Notice that the function \[ \sin x+\cos x+\tan x+\cot x+\sec x+\csc x \] is positive when\[ 0\le x\le \frac \pi 2 \]

OpenStudy (eliesaab):

the inequalities are strict above sorry

OpenStudy (eliesaab):

Also the above sum is negative when \[ \frac \pi2< x < 2 \pi \]

OpenStudy (eliesaab):

Knowing that you can split the interval into 2 intervals and get rid of the absolute values

OpenStudy (eliesaab):

The minimum in \( ] 0, 2\pi[\) occurs at \[ x=2 \pi +\tan ^{-1}\left(\frac{1-\sqrt{2}-\sqrt{2 \sqrt{2}-1}}{1-\sqrt{2}+\sqrt{2 \sqrt{2}-1}}\right) \]

Parth (parthkohli):

OK, I got it

Parth (parthkohli):